OpenStudy (bohotness):

Solve for y: y/5 > -5 y < -28 y < -25 y > -25 y > -28

OpenStudy (anonymous):

okayy umer

OpenStudy (anonymous):

multiply both sides by 5

OpenStudy (anonymous):

try every number for y and see which makes sense

OpenStudy (anonymous):

(y/5)*5 > -5*5 ...

OpenStudy (bohotness):

c

OpenStudy (anonymous):

that's correct

OpenStudy (amorfide):

\[\frac{ y }{ 5 }>-5\] since we have one fifth of y, we want to get a whole y, so how many fifths make 1? that is right! 5! so you multiply \[\frac{ y }{ 5}\] by 5 but since you did that to the left hand side you must also multiply the right hand side by 5 to make sure it is still equal this gives \[\frac{ 5y }{ 5 }>-5(5)\] this gives y>-25

OpenStudy (anonymous):

yes they are correct u may want to close this for another